The temperature has been as high as 89 and low as 43. Find values for a and b so that the solution to |x-a|_< b expresses the temperatures in terms of x. (The phone won't allow me to input less than or equal to, but that's what comes after the absolute value.
Answer by stanbon(75887)

The temperature has been as high as 89 and low as 43. Find values for a and b so that the solution to |x-a|_< b expresses the temperatures in terms of x.
-------------------------
Find the midpoint of the interval.
(89+43)/2 = 66
----
Find the distance to an endpoint.
89-66 = 23
-----
|x-66| <= 23
=================
Ans: a = 66 ; b = 23
==================
Cheers,
Stan H.
===============